What is the discount yield, bond equivalent yield, and effective annual return on a $5 million commercial paper issue that currently sells at 98.75 percent of its face value and is 131 days from maturity? (Use 360 days for discount yield and 365 days in a year for bond equivalent yield and effective annual return. Do not round intermediate calculations. Round your answers to 3 decimal places. (e.g., 32.161))
Answer #1

discount yield=(par value - purchase price)[/par value] * 360/days to maturity=(100-98.75)/100*360/131=3.435%

BEY=365*discount yield/(360-days to maturity*discount yield)=365*(100-98.75)/100*360/131/(360-131*(100-98.75)/100*360/131)=3.527%

EAR=(par/price)^(365/days to maturity)-1=(100/98.75)^(365/131)-1=3.567%